A firm is considering employing one of the two machines A and B over a period of 4 years at the end of which the salvage value of each is zero. The cost of machine A is $10, 000 while that of machine B is $11, 000. The probability distributions of the returns for each machine are given in the table below.
PROBABILTY MACHINE A($) MACHINE B($)
0.25 6000 7000
0.50 5000 5000
0.25 4000 3000
The risk free discount rate is 10% while the risk premium applied as follows.
STANDARD DEVIATION($) RISK PREMIUM
0 – 999 0%
1000 – 1999 10%
2000 – 2999 10%
3000 – 3999 20%

Which of the two machines should be installed?
1
Expert's answer
2018-01-16T09:20:07-0500
Machine A($) = (0.25 * 6000 + 0.5 * 5000+ 0.25 * 4000)/1.1^4 = 3415.07
Machine B($) = (0.25 * 7000 + 0.5 * 5000 + 0.25 * 3000)/1.1^4 = 3415.07
We get same results for both Machines, that`s why we can be installed or first Machine A, or Machine B.